The square of a number is 10 less than 7 times the number. Find the number.
Answer by mananth(16949)


Let the number be x
x^2=7x-10
x^2-7x+10=0
x^2-5x-2x+10=0
x(x-5)-2(x-5)=0
(x-5)(x-2)=0
x= 5 OR 2